OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Part 2: Provide the solution to the system. 2x – y = 5 x + 2y = 15

OpenStudy (anonymous):

Solve the first equation for y and substitute that into the seocnd equation

jhonyy9 (jhonyy9):

part 2. y=2x-5 and substitute into the second x+2(2x-5)=15 x+4x-10=15 5x=25 x=25/5 =5 x=5 y=2(5)-5=10-5=5 x=5 y=5